### v7

@@
Partial Identification of Individual-Level Parameters Using Aggregate Data in a Nonparametric Modelthanks: I am especially grateful to Edward Vytlacil, Isaiah Andrews, and Anna Mikusheva for their guidance and advice. I thank Alberto Abadie, Hyungsik Roger Moon, Whitney Newey, Jesse Shapiro, Elie Tamer, Vod Vilfort, and participants in the MIT econometrics lunch seminar for their helpful comments. This paper is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2141064. All errors are my own.
@@
I develop a methodology to partially identify linear combinations of conditional mean outcomes when the researcher only has access to aggregate data. Unlike the existing literature, I only allow for marginal, not joint, distributions of covariates in my model of aggregate data. Bounds are obtained by solving an optimization program and can easily accommodate additional polyhedral shape restrictions. I provide an empirical illustration of the method to Rhode Island standardized exam data.
@@ 2 Identification
One such restriction of interest is a polyhedral shape restriction on the conditional expectation function over groups $\mathbb{E}[Y_{i}|X_{i},G_{i}]$. This can be expressed as $S_{g}Y_{X,g}\leq a_{g}$ for each $g=1,\dots,G$, where $Y_{X,g}\equiv\left(\mathbb{E}[Y_{i}|X_{i}=x_{1},G_{i}=g],\dots,\mathbb{E}[Y_{i}|X_{i}=x_{K},G_{i}=g]\right)^{\prime}$, $S_{g}\in\mathbb{R}^{s_{g}\times K}$ are known fixed matrices, and $a_{g}$ are known fixed vectors. Another restriction of interest is a homogeneity restriction on the conditional average outcome over groups, which can be expressed as $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]=\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g^{\prime}]$ for all $k$ and $g,g^{\prime}$. It may also be the case that additional data at a finer level of aggregation is available to the researcher. For example, the researcher may also observe average group outcomes conditional on one of the $L$ covariates, $\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$. This additional data implies the restriction
$\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}\mathbb{E}[Y_{i}|X_{i}=x_{j},G_{i}=g]\mathbb{P}[X_{i}=x_{j}|G_{i}=g].$
@@ 4 Simulations
I present the true population aggregate features of the data for the first simulation study in Table 1, for the second simulation study in Table 2, and for the third simulation study in Table 3. In each study $\mathbb{E}[Y_{i}|G_{i}]$ is implied by the choice of marginal distributions for covariates $X_{i}$ and the above model. The data-generating process of the first study is calibrated to approximate the data used in the empirical illustration. In the second study, I maintain a similar marginal distribution for $X_{3i}$ but push the distributions for $X_{1i}$ and $X_{2i}$ much closer to 1 and 0 respectively. In the third study I maintain the same marginal distribution for $X_{2i}$ and $X_{3i}$, but instead enforce that some groups have almost all individuals with $X_{1i}=0$ and other groups have almost all individuals with $X_{1i}=1$.
@@
$G_{i}$ $\mathbb{E}[Y_{i}|G_{i}]$ $\mathbb{P}[X_{1i}=1|G_{i}]$ $\mathbb{P}[X_{2i}=1|G_{i}]$ $\mathbb{P}[X_{3i}=1|G_{i}]$ $\mathbb{E}[G_{i}=g]$ $\mathbb{P}[G_{i}=g]$
1 0.246 0.389 0.447 0.096 0.625 0.625
2 0.560 0.801 0.273 0.087 0.094 0.094
3 0.610 0.811 0.233 0.064 0.094 0.094
4 0.729 0.859 0.149 0.031 0.094 0.094
5 0.346 0.900 0.620 0.004 0.094 0.094
@@
$G_{i}$ $\mathbb{E}[Y_{i}|G_{i}]$ $\mathbb{P}[X_{1i}=1|G_{i}]$ $\mathbb{P}[X_{2i}=1|G_{i}]$ $\mathbb{P}[X_{3i}=1|G_{i}]$ $\mathbb{E}[G_{i}=g]$ $\mathbb{P}[G_{i}=g]$
1 0.553 0.732 0.194 0.096 0.625 0.625
2 0.712 0.893 0.126 0.087 0.094 0.094
3 0.744 0.896 0.111 0.064 0.094 0.094
4 0.810 0.914 0.079 0.031 0.094 0.094
5 0.744 0.931 0.189 0.004 0.094 0.094
@@
$G_{i}$ $\mathbb{E}[Y_{i}|G_{i}]$ $\mathbb{P}[X_{1i}=1|G_{i}]$ $\mathbb{P}[X_{2i}=1|G_{i}]$ $\mathbb{P}[X_{3i}=1|G_{i}]$ $\mathbb{E}[G_{i}=g]$ $\mathbb{P}[G_{i}=g]$
1 0.122 0.010 0.194 0.096 0.625 0.625
2 0.187 0.092 0.126 0.087 0.094 0.094
3 0.254 0.179 0.111 0.064 0.094 0.094
4 0.842 0.958 0.079 0.031 0.094 0.094
5 0.774 0.977 0.189 0.004 0.094 0.094
@@ 5 Empirical Illustration
In addition to the aggregate data as in Assumption 1.iv, the Rhode Island Department of Education also makes available pass rates conditional on each of the three covariates for school districts with a sufficiently high number of students in each subgroup. I use the following three subgroup pass rates: $\mathbb{E}[pass_{i}|white_{i}=1,G_{i}=g],\mathbb{E}[pass_{i}|econ_{i}=0,G_{i}=g]$, and $\mathbb{E}[pass_{i}|ELL_{i}=0,G_{i}=g]$. I drop all districts that are missing any of these subgroup pass rates before aggregating to the county level. The subgroup pass rates imply additional restrictions as discussed in Remark 2.
@@
I first present estimated bounds on math exam white/non-white pass rate gaps in Table 8 and on English exam white/non-white pass rate gaps in Table 9. 90% confidence intervals constructed as in Section 3.2 are displayed below the bound estimates in curly brackets. For both types of exam, bounds on the white/non-white pass rate gaps reported in Column 1 are wide and either uninformative or close to uninformative. Imposing the additional monotonicity restrictions helps narrow the bounds for some parameters, reported in Column 2, but bounds are still wide. Using additional subgroup pass rates without monotonicity in Column 3 further narrows the bounds for those same parameters. Adding monotonicity to additional subgroup pass rates in Column 4 produces the tightest bounds, especially for white/non-white pass rate gaps among students who are economically disadvantaged but not English-language learners. For example, among economically disadvantaged students who are not English-language learners, the white/non-white English exam pass rate gap is estimated to be no lower than -30% and no higher than 64%.
@@
In Table 10 I present estimated bounds on math exam conditional pass rates and in Table 11 I present estimated bounds on English exam conditional pass rates. 90% confidence intervals constructed as in Section 3.2 are displayed below the bound estimates in curly brackets. Just as for the white/non-white pass rate gaps, the additional restrictions help to narrow the bounds. Bounds on certain conditional pass rates are informative with additional subgroup pass rates under monotonicity. For example, the math exam pass rate among non-white students who are economically disadvantaged and not English-language learners is estimated to be bounded above by 29%, while the math exam pass rate among white students who are not economically disadvantaged and not English-language learners is estimated to be between 37% and 72%. Similarly, the English exam pass rate among non-white students who are economically disadvantaged and not English-language learners is estimated to be bounded above by 36%, while the math exam pass rate among white students who are not economically disadvantaged and not English-language learners is estimated to be between 50% and 80%.
@@ A.1 Bounds using the Cross and Manski (2002) method
Cross and Manski (2002) and Cho and Manski (2008) provide sharp bounds on $\mathbb{E}[Y_{i}|X_{i},G_{i}]$ jointly over the support of $X_{i}$ when I know the probability distribution of $Y_{i}|G_{i}$ and the probability distribution of $X_{i}|G_{i}$. In my setting instead of the distribution of $Y_{i}|G_{i}$ I observe $\mathbb{E}[Y_{i}|G_{i}]$ and instead of the joint distribution of $X_{i}|G_{i}$ I observe the marginal distribution of each $X_{\ell i}|G_{i}$.

### v8

@@
Partial Identification of Individual-Level Parameters Using Aggregate Data in a Nonparametric Model
@@
I develop a methodology to partially identify linear combinations of conditional mean outcomes when the researcher only has access to aggregate data. Unlike the existing literature, I only allow for marginal, not joint, distributions of covariates in my model of aggregate data. Bounds are obtained by solving an optimization program and can easily accommodate additional polyhedral shape restrictions. I provide a procedure to construct confidence intervals on the identified set and demonstrate performance of my method in a simulation study. In an empirical illustration of the method using Rhode Island standardized exam data, I find that conditional pass rates vary across student subgroups and across counties.
@@ 2 Identification
One such restriction of interest is a polyhedral shape restriction on the conditional expectation function over groups. This can be expressed as $S_{g}c_{g}\leq a_{g}$ for each $g=1,\dots,G$, where $c_{g}\equiv\left(c_{1g},\dots,c_{Kg}\right)^{\prime}$, $S_{g}\in\mathbb{R}^{s_{g}\times K}$ are known fixed matrices, and $a_{g}$ are known fixed vectors. Another restriction of interest is a homogeneity restriction on the conditional average outcome over groups, which can be expressed as $c_{kg}=c_{kg^{\prime}}$ for all $k$ and $g,g^{\prime}$. It may also be the case that additional data at a finer level of aggregation is available to the researcher. For example, the researcher may also observe average group outcomes conditional on one of the $L$ covariates, $\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$. This additional data implies the restriction
$\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{jg}p_{jg}.$
@@ 4 Simulations
I present the true population aggregate features of the data for the first simulation study in Table 1, for the second simulation study in Table 2, and for the third simulation study in Table 3. In each study $\mathbb{E}[Y_{i}|G_{i}]$ is implied by the choice of marginal distributions for covariates $X_{i}$ and the above model. The data-generating process of the first study is calibrated to approximate the data used in the empirical illustration. In the second study, I maintain the same marginal distribution for $X_{3i}$ but push the distributions for $X_{1i}$ and $X_{2i}$ much closer to 1 and 0 respectively. In the third study I maintain the same marginal distribution for $X_{2i}$ and $X_{3i}$ as in the second study, but instead enforce that some groups have almost all individuals with $X_{1i}=0$ and other groups have almost all individuals with $X_{1i}=1$.
@@
$G_{i}$ $\mathbb{E}[Y_{i}|G_{i}]$ $\mathbb{P}[X_{1i}=1|G_{i}]$ $\mathbb{P}[X_{2i}=1|G_{i}]$ $\mathbb{P}[X_{3i}=1|G_{i}]$ $\mathbb{P}[G_{i}=g]$
1 0.246 0.389 0.447 0.096 0.625
2 0.560 0.801 0.273 0.087 0.094
3 0.610 0.811 0.233 0.064 0.094
4 0.729 0.859 0.149 0.031 0.094
5 0.346 0.900 0.620 0.004 0.094
@@
$G_{i}$ $\mathbb{E}[Y_{i}|G_{i}]$ $\mathbb{P}[X_{1i}=1|G_{i}]$ $\mathbb{P}[X_{2i}=1|G_{i}]$ $\mathbb{P}[X_{3i}=1|G_{i}]$ $\mathbb{P}[G_{i}=g]$
1 0.553 0.732 0.194 0.096 0.625
2 0.712 0.893 0.126 0.087 0.094
3 0.744 0.896 0.111 0.064 0.094
4 0.810 0.914 0.079 0.031 0.094
5 0.744 0.931 0.189 0.004 0.094
@@
$G_{i}$ $\mathbb{E}[Y_{i}|G_{i}]$ $\mathbb{P}[X_{1i}=1|G_{i}]$ $\mathbb{P}[X_{2i}=1|G_{i}]$ $\mathbb{P}[X_{3i}=1|G_{i}]$ $\mathbb{P}[G_{i}=g]$
1 0.122 0.010 0.194 0.096 0.625
2 0.187 0.092 0.126 0.087 0.094
3 0.254 0.179 0.111 0.064 0.094
4 0.842 0.958 0.079 0.031 0.094
5 0.774 0.977 0.189 0.004 0.094
@@ 5 Empirical Illustration
In addition to the aggregate data of Assumption 1.iv, the Rhode Island Department of Education also makes available pass rates conditional on each of the three covariates for school districts with a sufficiently high number of students in each subgroup. I use the following three subgroup pass rates: $\mathbb{E}[pass_{i}|white_{i}=1,G_{i}=g],\mathbb{E}[pass_{i}|econ_{i}=0,G_{i}=g]$, and $\mathbb{E}[pass_{i}|ELL_{i}=0,G_{i}=g]$. I drop all districts that are missing any of these subgroup pass rates before aggregating to the county level. The subgroup pass rates imply additional restrictions as discussed in Remark 2.
@@
I first present estimated bounds on math exam white/non-white pass rate gaps in Table 8 and on English exam white/non-white pass rate gaps in Table 9. 90% confidence intervals constructed as in Section 3.2 are displayed below the bound estimates in curly brackets. For both types of exam, bounds on the white/non-white pass rate gaps reported in Column 1 are wide and either uninformative or close to uninformative. Imposing the additional monotonicity restrictions helps narrow the bounds for some parameters, reported in Column 2, but bounds are still wide. Using additional subgroup pass rates without monotonicity in Column 3 further narrows the bounds for those same parameters. Adding monotonicity to additional subgroup pass rates in Column 4 produces the tightest bounds, especially for white/non-white pass rate gaps among students who are economically disadvantaged but not English-language learners. For example, among economically disadvantaged students who are not English-language learners, the white/non-white English exam pass rate gap is estimated to be no lower than -30% and no higher than 64%. Bounds on pass rate gaps among English-language learners are never informative, as expected due to the small percentage of English-language learners in Rhode Island.
@@
In Table 10 I present estimated bounds on math exam conditional pass rates and in Table 11 I present estimated bounds on English exam conditional pass rates. 90% confidence intervals constructed as in Section 3.2 are displayed below the bound estimates in curly brackets. Just as for the white/non-white pass rate gaps, the additional restrictions help to narrow the bounds on pass rates for students who are not English-language learners. For example, the math exam pass rate among non-white students who are economically disadvantaged and not English-language learners is estimated to be bounded above by 29%, while the math exam pass rate among white students who are not economically disadvantaged and not English-language learners is estimated to be between 37% and 72%. Similarly, the English exam pass rate among non-white students who are economically disadvantaged and not English-language learners is estimated to be bounded above by 36%, while the math exam pass rate among white students who are not economically disadvantaged and not English-language learners is estimated to be between 50% and 80%.
@@ 6 Conclusion
Acknowledgements
I am especially grateful to Edward Vytlacil, Isaiah Andrews, and Anna Mikusheva for their guidance and advice. I thank Editor Yuya Sasaki, the Associate Editor, two anonymous referees, Alberto Abadie, Hyungsik Roger Moon, Whitney Newey, Jesse Shapiro, Elie Tamer, Vod Vilfort, and participants in the MIT econometrics lunch seminar for their helpful comments and suggestions. All errors are my own.
Funding
This paper is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2141064.
Disclosure Statement
There are no competing interests to declare.
@@ A.1 Bounds using the Cross and Manski (2002) method
Cross and Manski (2002) and Cho and Manski (2008) provide sharp bounds on $\mathbb{E}[Y_{i}|X_{i},G_{i}]$ jointly over the support of $X_{i}$ when one knows the probability distribution of $Y_{i}|G_{i}$ and the probability distribution of $X_{i}|G_{i}$. In my setting instead of the distribution of $Y_{i}|G_{i}$ I observe $\mathbb{E}[Y_{i}|G_{i}]$ and instead of the joint distribution of $X_{i}|G_{i}$ I observe the marginal distribution of each $X_{\ell i}|G_{i}$.